OpenStudy (anonymous):

Simplify the expression. Square root -108 + square root -27.

OpenStudy (anonymous):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

well their has to be a way is one of my questions.... heres the possible answers

OpenStudy (anonymous):

-3 square root 15 3i square root 15 9i square root 3 -9 square root 3

OpenStudy (anonymous):

can you use the equation button below to write the question

OpenStudy (anonymous):

\[\sqrt{-108} + \sqrt{-27}\]

OpenStudy (anonymous):

\[\sqrt{-108}=\sqrt{-36*3}\] can you finish it

OpenStudy (anonymous):

\[\sqrt{-27}=\sqrt{-9*3}\] ??

OpenStudy (anonymous):

yes now how can you rewrite it

OpenStudy (anonymous):

-9 square root 3?

OpenStudy (anonymous):

square root of -9 = 3i

OpenStudy (anonymous):

i know but would that be the answer ti my question?